Ask
your own question, for FREE!
Mathematics
11 Online
A basket contains 11 pieces of fruit: 4 apples, 5 oranges, and 2 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple?
Still Need Help?
Join the QuestionCove community and study together with friends!
9/10?
Total outcomes = 11 favorable outcomes = 5 Probability = Number of favorable outcomes / Total number of outcomes Probability that Jonas will get an orange = 5/11 Total outcomes = 10 (because 1 fruit is already taken by Jonas) favorable outcomes = 4 Probability Beth will get an apple = 4/10 = 2/5
so 9/11?
no , multiply both probability that Jonas will get an orange and Beth will get an apple = 5/11 * 2/5 = 2/11
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
6 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal